# Hamiltonian simulation-based quantum-selected configuration interaction for large-scale electronic structure calculations with a quantum computer

### v1

@@ V Conclusion
In this study, we proposed an HSB-QSCI, which uses a quantum computer to perform Hamiltonian simulation and subsequent measurement of the quantum state in the computational basis to sample the important Slater determinants used to span the subspace Hamiltonian to be diagonalized on a classical computer. Compared to the reported state preparation approaches for the QSCI, the HSB-QSCI is free from variational optimizations of the quantum states, and it works with the simple initial wave function like the HF. Since long-time evolution is in general difficult to simulate on a classical computer, we expect this approach to become powerful as the system size increases. Proof-of-concept simulations were performed for the lowest spin-singlet and spin-triplet states of organic molecules including benzene, naphthalene, anthracene, phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene, and we achieved chemical precision in all molecules studied. The quantum hardware demonstrations of the HSB-QSCI for the lowest spin-singlet and spin-triplet states of hexa-1,2,3,4,5-pentaene were also reported using ibm_torino processor with 20 qubits, with the aid of MPO-based classical optimization of the quantum circuit for the time evolution operator. By setting the number of shots to and using the SCCR, we obtained the HSB-QSCI total energies with a deviation from the CAS-CI values of at most 0.15 kcal mol-1, illustrating the usefulness of the HSB-QSCI approach.
@@ VI Acknowledgments
S1.4 Shot number dependence on the HSB-QSCI simulations of anthracene
S1.5 Hardware results of the HSB-QSCI with three-configurational wave functions

### v2

@@ V Conclusion
In this study, we proposed an HSB-QSCI, which uses a quantum computer to perform Hamiltonian simulation and subsequent measurement of the quantum state in the computational basis to sample the important Slater determinants used to span the subspace Hamiltonian to be diagonalized on a classical computer. Compared to the reported VQE-based state preparation approaches for the QSCI [14, 16, 17], the HSB-QSCI is free from variational optimizations of the quantum states, and it works with the simple initial wave function like the HF. Since long-time evolution is in general difficult to simulate on a classical computer, we expect this approach to become powerful as the system size increases. Proof-of-concept simulations were performed for the S0 and T1 states of organic molecules 1–5, and we achieved chemical precision in all molecules studied. The quantum hardware demonstrations of the HSB-QSCI for the S0 and T1 states of three carbyne molecules 5–7 were also reported using ibm_kawasaki with 20, 28, and 36 qubits, respectively, with the aid of MPO-based classical optimization of the quantum circuit for the time evolution operator. The HSB-QSCI achieved chemical precision for the total energy in 5, for the singlet–triplet energy difference in 6, and for the energy difference between planar and twisted geometries of the S0 state in 7. In 7, by considering about 1% of the Slater determinants, the HSB-QSCI captured more than 99.75% of correlation energies in the active space. The largest system studied in this work is 7 with 36 qubits, which is the largest system capable of performing the CAS-CI calculation in a typical computational environment without a supercomputer [45, 55]. We expect that the HSB-QSCI can handle the system larger than 36 qubits, by introducing batch-based implementations in the Hamiltonian matrix diagonalization part, and calculating the variance of the energy expectation value as explored in Reference [19]. Note that, as discussed by Reinholdt and coworkers in their recent paper [56], the QSCI wave function is in some cases less compact than that obtained by classical heuristics. From Eq. (6), the probability that a particular Slater determinant is measured in the HSB-QSCI depends not only on the CI coefficient of the ground state, but also on the gap to the excited state and the CI coefficient of the excited state, and screening out unimportant Slater determinants from the subspace is very important in the application to larger systems. Applying the auxiliary-field quantum Monte Carlo (AFQMC) method with the QSCI wave function as the input can be another direction towards larger-scale quantum chemical calculations with a quantum computer, as suggested in recent studies [57, 58]. Application of the HSB-QSCI to larger systems is currently in progress.
@@ VI Acknowledgments
S1.4 HSB-QSCI results calculated using kernel_fixed_space subroutine in PySCF
S1.5 Shot number dependence on the HSB-QSCI simulations of 3
S1.6 Shot number dependence on the HSB-QSCI simulations of 5 with the maximum locality